OpenStudy (anonymous):

Could someone please assist me with this (piecewise function): Find the numbers at which f(x)= {4+x^2 x less than 0} {10+x 0

OpenStudy (anonymous):

Check the limits of the function as you reach each "endpoint" of each interval of each piece. In other words, for instance, to see whether \(f\) is discontinuous at \(x=0\), find \(\displaystyle\lim_{x\to0^-}f(x)\) and \(\displaystyle\lim_{x\to0^+}f(x)\). If these limits do not match (not equal) then \(f\) is discontinuous at \(x=0\). Also, these limits must be the same as \(f(0)\). For the first limit, \(x\to0\) from the left, which translates to "consider values of \(x\) to the left of \(x=0\)," or \(x<0\). This means you use the piece of \(f\) that is defined for this interval; namely \(f(x)=4+x^2\).